OpenStudy (anonymous):

multiply and simplify by factoring 4root 10,000 * 4root 100,000

OpenStudy (anonymous):

root of 10,000 is 100 x 10 which is 10 root(10) root of 100,000 is 100 x 100 which is 100 40root(10) + 400

OpenStudy (anonymous):

4 sqrt 10,000 * 4 sqrt 100,000

OpenStudy (anonymous):

\[\sqrt{10,000}= \sqrt{100\times10} = 10\sqrt{10}\] \[\sqrt{100,000}= \sqrt{100\times10} = 100\]

OpenStudy (anonymous):

\[40\sqrt{10} \times400 = 16,000\sqrt{10}\]

OpenStudy (anonymous):

4 on top of the sqrt sign 10,000

OpenStudy (anonymous):

yeah so that 4 is multiplied by 10 that we brought out of the square root sign to get 40sqrt(10)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do you get it?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

so what is the answer

OpenStudy (anonymous):

the answer is 16,000sqrt(10)

OpenStudy (anonymous):

got it thanks

OpenStudy (anonymous):

I have 7 more questions is it ok to post them?

OpenStudy (anonymous):

lol yeah go ahead...I'll do a few with you and make sure you get it so that you can do the rest.

OpenStudy (anonymous):

thanks :-)

OpenStudy (anonymous):

There's no limit to # questions you can post..you're welcome

OpenStudy (anonymous):

sounds good

OpenStudy (anonymous):

Add simplify by collecting like radical terms, if possible. Assume that all expressions under radicals represent non-negative numbers 4 sqrt y + 3 sqrt y =

OpenStudy (anonymous):

carla is it \[\sqrt[4]{10000}\] or \[4\sqrt{10000}\]?

OpenStudy (anonymous):

Okay..so before you can add any radical term you have to make sure that the expression in the radials are the same. In this case they are we have y in both square roots. So all we have to do is to add the coefficient of the square roots which are 4 and 3. \[4\sqrt{y} + 3\sqrt{y} = 7\sqrt{y}\]

OpenStudy (anonymous):

yeah satellite it keyboard error for my first solution so its not correct. Carla are you with me so far?